# SWAN: SGD with Normalization and Whitening Enables Stateless LLM Training

### v2

@@ Results
Figure 5 compares SWAN’s actual performance against the counterfactual additive curves. If SWAN’s actual curves exceed these additive estimates, it indicates a tendency towards a multiplicative speedup, instead of the additive advantage. We summarize the observations for model sizes of 60M, 130M, 350M, and 1.3B parameters:
  • For smaller models (60M and 130M), the actual speedup trajectories align closely with the additive baseline, indicating a primarily additive speedup.
  • For larger models (350M and 1.3B), the actual curves rise noticeably above the additive estimates, suggesting a multiplicative speedup. This indicates that SWAN yields increasing efficiency gains as model size grows.
6.3 Memory Efficiency and Throughput Analysis
In this section, we measure both the memory footprint and throughput of SWAN.
Memory Footprint
We compare SWAN, Adam, and Galore on a single A100 GPU. Unlike Zhao et al. (2024a), which reports layer-wise training memory usage, we measure total end-to-end memory consumption under full-model training using batch size of $1$ for LLaMA with 1.3B, 7B, and 13B parameters. As shown in Figure 1 (c), SWAN’s memory usage is close to that of vanilla SGD, providing nearly a 50% reduction in total memory by removing the needs for first and second moments compared to Adam. This observation underlines the benefit of a stateless design.
Effective Throughput
We also assess throughput when training a 1.3B-parameter model on 4 A100 GPUs with a token batch size of 130K tokens. We use two metrics:
  • Absolute throughput: number of tokens processed per second.
  • Effective throughput: absolute throughput adjusted by SWAN’s token efficiency relative to Adam.
These metrics evaluate whether $\mathtt{GradWhitening}$ (which uses Newton-Schulz iterations) significantly reduces training speed and also account for the fact that some optimizers make more effective use of training tokens. As shown in Figure 1 (d), SWAN achieves an absolute throughput comparable to Adam, indicating that the computational overhead of $\mathtt{GradWhitening}$ is minimal. Consequently, SWAN exhibits a higher effective throughput than Adam because of its more efficient use of tokens during training. All experiments use BF16 precision for $\mathtt{GradWhitening}$.
@@ 6.4 Ablation Studies
We consider six ablation settings: (1) SWAN(full), (2) SWAN ($\mathtt{GradNorm}$ only), (3) SWAN ($\mathtt{GradWhitening}$ only), (4) Adam (full), (5) Adam (momentum only), and (6) Adam (second moment only). As shown in Figure 6 (a), both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to SWAN’s final performance. Removing either results in performance degradation. Similarly, Adam also requires first and second moment estimates for optimal performance.
@@
6.5 Analysis of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
6.5.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
Baselines
We compare:
Methodology
Results
6.5.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
(a) Quadratic optimization, well-conditioned
(b) Quadratic optimization, ill-conditioned
(c) Rastrigin function optimization
(d) Effect of initialization
Baselines
Results
From Figure 8 (a)–(c), we summarizes the following outcomes:
@@ Appendix A Desired properties of adaptive optimizers
Appendix B Additional Analysis
B.1 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
@@ Appendix B Additional Analysis
B.2 Numerical Verification of Proposition 1
@@
Appendix C Proof of Theorem 1
@@
$\dot{\tilde{{\bm{u}}}}_{i}$ $=\frac{\alpha_{i}(t)(\mu_{i}+{\bm{\xi}}_{i})-\alpha_{i}(t)(\frac{1}{n}\sum_{j}\mu_{ij}+\frac{1}{n}\sum_{j}\xi_{ij})}{\alpha_{i}(t)(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij}-\frac{1}{n}\sum_{j}\mu_{ij}-\frac{1}{n}\sum_{j}\xi_{ij})^{2})}$
$=\frac{(\mu_{i}+{\bm{\xi}}_{i})-(\frac{1}{n}\sum_{j}\mu_{ij}+\frac{1}{n}\sum_{j}\xi_{ij})}{(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij}-\frac{1}{n}\sum_{j}\mu_{ij}-\frac{1}{n}\sum_{j}\xi_{ij})^{2})}$
@@
Appendix D Proof of Theorem 2
@@
Appendix E Proof of Proposition 2
@@
Appendix F Proof of Proposition 3
@@
Appendix G Proof of Proposition 1
@@
Appendix H Implementation details
Setup
@@ Appendix H Implementation details
We describe the implementation setups for SWAN used in LLM pre-training tasks. To enable a more straightforward and comparable analysis, we simply replicate the setting of Zhao et al. (2024a), under exactly the same model configs and optimizer configs. This includes the same model architecture, tokenizer, batch size, context length, learning rate scheduler, etc. The only exception is that since SWAN does not need warm-up, we use zero-warm-up ratio for SWAN. But we still ensure that the learning rates of all methods are fully decayed to the same target learning rate (0.1 $\times$ initial learning rate). Across all model sizes, we use the same hyperparameter settings and the same learning rate tuning procedure. The only exception is that for a model of size 1.3B: as we already know that a larger model requires smaller learning rates, we conduct a learning search for Adam over a smaller but more fine-grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$. As a result, the optimal learning rate found for Adam on 1.3B is 0.0007.
Applying SWAN on LLMs
@@
Since SWAN utilizes matrix-level operations on gradients, it can only be applied to 2D parameters. Therefore, in our experiments, we only apply $SWAN{}$ on all linear projection weights in transformer blocks. Similar to Galore (Zhao et al., 2024a), the rest of the non-linear parameters still uses Adam as the default choice. Therefore, we follow the learning rate setup of Galore, where we fix a global learning rate 0.001 across all model sizes and all modules. Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate. We choose $\alpha=1$ for all model sizes except for 1.3 B case. This is because we observe that the optimal learning rate of Adam under 1.3B becomes smaller than 0.001, hence we also reduce the learning rate on SWAN, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$. Note that this choice is quite arbitrary, and it is not extensively grid-searched. This is to show how the method work out-of-the box, without extensive problem-specific tuning.
Implementation of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
For $\mathtt{GradNorm}$, we simply set $\bar{{\bm{g}}}=0$ to simplify computation, as we found in practice this often has small scales. For $\mathtt{GradWhitening}$, before N-S iteration, we further normalize its input matrix by its Frobenius norm. This is applied in all our ablation studies as well, regardless of whether the gradient is processed by $\mathtt{GradNorm}$. For the N-S iteration, we found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence. We set $\beta=0.8$ with iteration number $k=10$ and it generally gives satisfactory performance. $\mathtt{GradWhitening}$ can be performed in both FP32 and BF16 format. In FP32 format, we first convert the gradient matrice to FP32, perform $\mathtt{GradWhitening}$, and then convert back to BF16. This is the default setup in Section 6.1 unless stated otherwise. Alternatively, $\mathtt{GradWhitening}$ can also be done in BF16 precision. The setup we use for BF16 precision is $\beta=0.85$, $k=6$. These hyperparameters are selected on randomly sampled matrices via minimizing inverse square root estimation loss; then directly applied to LLM without any further tuning. We found that this usually translates well. Note that this might be further optimized similar to (Jordan et al., 2024), we leave it to future work as N-S performance optimization is orthogonal to our key contribution.

### v3

@@ Results
Figure 5 compares SWAN’s actual performance against the counterfactual additive curves. If SWAN’s actual curves exceed these additive estimates, it indicates a tendency towards a multiplicative speedup, instead of the additive advantage. We summarize the observations across model sizes as: 1) for smaller models (60M and 130M), the actual speedup trajectories align closely with the additive baseline, indicating a primarily additive speedup; 2) for larger models (350M and 1.3B), the actual curves rise noticeably above the additive estimates, suggesting a multiplicative speedup. This indicates that SWAN yields increasing efficiency gains as model size grows.
@@ 6.4 Ablation Studies
We take SWAN-0 from the SWAN series as an example and conduct the following ablation studies. All SWAN optimizer mentioned in this section specifically refers to the SWAN-0 setting.
@@
We consider six ablation settings: (1) SWAN(full), (2) SWAN ($\mathtt{GradNorm}$ only), (3) SWAN ($\mathtt{GradWhitening}$ only), (4) Adam (full), (5) Adam (momentum only), and (6) Adam (second moment only). As shown in Figure 6 (a), both $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ contribute to SWAN’s final performance. Removing either results in performance degradation. Similarly, Adam also requires all moments for optimal performance.
@@
7 Open Questions and Extensions
In Section 5 we analyzed the design choice of $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$ from the perspective of LLM dynamics. However, the compositional effect of these two operators is still an open question, which is not discussed in this work. Specifically, in Appendix B we empirically show that when comparing with standard, standalone NS processing of the gradients, both $\mathtt{GradWhitening}$ and (NSDS)-$\mathtt{GradWhitening}$ of SWAN introduces significant changes, rotating the update by a relatively large magnitude. This indicates that additional efforts are needed to further explain the novel dynamics of SWAN, as well as the question of “what makes a good preprocessing chain for gradients”.
After the v1 version of this paper on arXiv, we posted a follow-up work of SWAN (Scetbon et al., 2025), in which we have shown that SWAN is a special case of a general framework called multi-normalized gradient descent (MNGD). MNGD aims at normalizing gradients according to multiple norms, generalizing the steepest descent viewpoint of Bernstein & Newhouse (2024b) that recasts popular first-order optimizers as normalization of gradients under a single norm.
@@ Appendix A Desired properties of adaptive optimizers
Appendix B Acceleration of Newton-Schulz iteration via diagonal substitution
B.1 Algorithnms
Computing $\mathtt{GradWhitening}$ exactly can be expensive, as it involves solving the matrix square-root inverse. One option is to directly apply the Newton-Schulz variant of decorrelated batch normalization (Song et al., 2022; Li et al., 2018; Huang et al., 2019), which allows a more GPU-friendly estimation. This is given by (Song et al., 2022; Li et al., 2018):
$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k})\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-{\mathbf{Z}}_{k}{\mathbf{Y}}_{k}){\mathbf{Z}}_{k}\end{cases}$
where ${\mathbf{Y}}_{0}={\mathbf{G}}{\mathbf{G}}^{\top}$ , ${\mathbf{Z}}_{0}={\mathbf{I}}$. At convergence, $\mathtt{GradWhitening}({\mathbf{G}})={\mathbf{Z}}{\mathbf{G}}$ (Algorithm 2). However, estimating $({\mathbf{G}}{\mathbf{G}}^{\top})^{-1/2}$ with NS requires $\mathcal{O}(m^{3})$ (assuming $m<n$) complexity. Here, we propose a heuristic scheme that has $\mathcal{O}(m^{2})$ complexity to estimate square-root inverse:
$\begin{cases}{\mathbf{Y}}_{k+1}=\frac{1}{2}{\mathbf{Y}}_{k}\text{Diag}(3{\mathbf{I}}-{\mathbf{Z}}_{k}\text{Diag}({\mathbf{Y}}_{k})))\\ {\mathbf{Z}}_{k+1}=\frac{1}{2}(3{\mathbf{I}}-\text{Diag}({\mathbf{Z}}_{k}){\mathbf{Y}}_{k})\text{Diag}({\mathbf{Z}}_{k})\end{cases}$
where $\text{Diag}(\cdot)$ returns a diagonal matrix that has the same diagonal elements as the input matrix. Basically, whenever we encounter matrix multiplication in NS iterations, we replace one of them by its diagonal approximation. We refer to this as the NS with diagonal substitution (NSDS) scheme.
Note that in the above standard presentation we have fixed that both NS and NSDS uses coefficients $=0.5$ for ${\mathbf{Y}}$ updates and ${\mathbf{Z}}$ updates, respectively. In practice we may further tune these coefficients to compensate for short number of iterations (usually under 10).
B.2 Experiment: LLM Gradient Condition Number Reduction
Setup
In this synthetic experiment, we assessed the effectiveness of two whitening methods, Newton-Schulz and the proposed Newton-Schulz with diagonal substitution (NSDS) scheme, on gradient matrices obtained from LLM training. As discussed in Section 4.1, the exact $\mathtt{GradWhitening}$ operator results in matrices with optimal condition number ($=1$); therefore, we hereby investigate the matrix condition numbers of the processed gradients obtained from different methods. Specifically, both methods use 5 NS iterations with NS step size optimized. We train a 130M LLama model following the architecture setting of Section 6.1 on randomly generated sequences for 1000 steps, and take the MLP weights of a middle layer (we take the fifth layer without loss of generality) and use different methods to whiten the gradient matrices. We consider three methods: standard NS; NSDS; and SWAN with NSDS (that is, composing $\mathtt{GradNorm}$ with NSDS-$\mathtt{GradWhitening}$). At each training step, the condition number reduction ratio of different method was calculated for both whitening methods (the higher the better). Note that for all methods, the gradients have been pre-normalized by its norm.
Results
Results are shown in Figure 7 (a). We notice that NSDS alone (orange curve) is not sufficient to reach a good condition number reduction ratio. However, when combined with $\mathtt{GradNorm}$ (i.e., SWAN with NSDS, the green curve), its performance started to catch up and even outperform the standard NS method after 500 training steps. This show the effectiveness of the proposed scheme. One potential caveat that we spot is that the condition number produced by SWAN with NSDS is more noisy than the standard NS iteration; which might lead to improvements that will be addressed in future work.
The significance of $\mathtt{GradNorm}$ and NSDS
[…]
Figure 7: Comparison between different whitening schemes. (a) Performance comparison by condition number reduction ratio (higher the better). (b) Cosine similarities between the whitened matrix produced by different preprocessing schemes, respectively.
B.3 Ablation: effect of NSDS iterations
We examine the impact of the number of iterations of our Newton Schulz with Diagonal Substitution (NSDS) scheme. With SWAN, we compare the test PPL performance on 130M model. Results are shown in Table 4. As we can see, the improvement brought by additional NSDS iteration is marginal; hence in this paper, we only apply 2 iterations of NSDS.
Table 4: Effect of NSDS iterations on test PPL. Results are obtain using FP32 precision.
# NSDS iterations Test PPL
1 - (LLM loss diverge)
2 22.63
5 22.62
10 22.61
B.4 Ablation: Precision
We compare the performance of different precisions of $\mathtt{GradWhitening}$ with NSDS scheme. As shown in Table 5, on 130M model (2 step NSDS iterations) ablation we show that BF16 can be used without major performance degrade compared to FP32.
Table 5: BF16 vs FP 32 NSDS
NSDS Precision Test PPL on 130M
BF16 22.61
FP32 22.63
Appendix C Additional Analysis
C.1 Analyzing the GradWhitening Pt. II: Robustness Against Local Curvature
@@ Appendix C Additional Analysis
C.2 Numerical Verification of Proposition 1
@@
Appendix D Proof of Theorem 1
@@
$\dot{\tilde{{\bm{u}}}}_{i}$ $=\frac{\alpha_{i}(t)(\mu_{i}+{\bm{\xi}}_{i})}{\alpha_{i}(t)(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij})^{2})}$
$=\frac{(\mu_{i}+{\bm{\xi}}_{i})}{(\frac{1}{n}\sum_{j}(\mu_{ij}+\xi_{ij})^{2})}$
@@
Appendix E Proof of Theorem 2
@@
Appendix F Proof of Proposition 2
@@
Appendix G Proof of Proposition 3
@@
Appendix H Proof of Proposition 1
@@
Appendix I Empirical verification of theoretical insights from Section 5 regarding $\mathtt{GradNorm}$ and $\mathtt{GradWhitening}$
I.0.1 Does $\mathtt{GradNorm}$ Stabilize Gradient Distributions of SGD?
Baselines
We compare:
Methodology
Results
I.0.2 Does $\mathtt{GradWhitening}$ Counteracts Local Curvature and Provide Fast Convergence on ill-conditioned problems?
(a) Quadratic optimization, well-conditioned
(b) Quadratic optimization, ill-conditioned
(c) Rastrigin function optimization
(d) Effect of initialization
Baselines
Results
From Figure 10 (a)–(c), we summarizes the following outcomes:
Appendix J Implementation details
General setup
@@ Results
We describe the implementation setups for SWAN used in LLM pre-training tasks. To enable a more straightforward and comparable analysis, we simply replicate the setting of Zhao et al. (2024a), under exactly the same model configs and optimizer hyperparameter configs, whenever possible. This includes the same model architecture, tokenizer, batch size, context length, learning rate scheduler, learning rates, subspace scaling, etc.
Precision
[…]
Learning rate scheduling
we use exactly the same scheduler as in Zhao et al. (2024a), with the exception of SWAN-0, which does not require any learning rate warmup. Therefore, for SWAN-0, we directly start with maximum learning rate, and enter the learning rate decay phase, using the same decay parameters as Zhao et al. (2024a).
Reproducing baseline results
Most baseline results are cited from respective papers (Zhao et al., 2024a; Zhu et al., 2024) as we share the exact same setup. We also tried to reproduce their results using the same opensourced code, and generally obtain slightly worse results for Galore, Apollo and Adam for larger models (350M and 1B, see Table 6). Therefore in the main paper we only compare with the official results. For the reproduced results of Adam, we specify the details below as it was not disclosed in Zhao et al. (2024a). We use same learning rate tuning procedure as suggested by Zhao et al. (2024a) (i.e., performing grid search over $\{0.01,0.005,0.001,0.0005,0.0001\}$). We found that the optimal learning rates for Adam is 0.001. The only exception is that for a model of size 1.3B: as we already know that a larger model requires smaller learning rates, we conduct a learning search for Adam over a smaller but more fine-grained grid of $\{0.001,0.0007,0.0005,0.0003,0.0001\}$. As a result, the optimal learning rate found for Adam on 1.3B is 0.0007. Finally, one baseline that does not exist in the literature is the Momentum + $\mathtt{GradWhitening}$ (Muon-like optimizer without Nestrov acceleration) baseline; and we report our own results. We start from the default learning rates used by Muon Jordan et al. (2024) and tuned them over a grid of ${0.01,0.02,0.03,0.04,0.05}$.
Table 6: Reproduced results.
130M 350M 1.3 B
Adam (reproduced) 24.44 (0.75G) 19.24 (2.05G) 16.44 (7.48G)
Apollo-mini (reproduced) 23.97 (0.43G) 17.60 (0.93G) 14.37 (2.98G)
Galore (reproduced) 24.67 (0.57G) 19.74 (1.29G) 15.89 (4.43G)
$r$ of low-rank methods 256 256 512
Training Steps 20K 60K 100K
SWAN Settings and hyperparameters
@@
Since SWAN utilizes matrix-level operations on gradients, it can only be applied to 2D parameters. Therefore, in our experiments, we only apply SWAN on all linear projection weights in transformer blocks. Similar to Galore (Zhao et al., 2024a), the rest of the non-linear parameters still uses Adam as the default choice. Therefore, we follow the learning rate setup of Galore, where we fix some global learning rate across all model sizes and all modules. Then, for the linear projection modules where SWAN is applied, we simply apply a scaling factor $\alpha$ on top of the global learning rate. For all SWAN variants, we adopt a lazy-tuning approach (hyperparameters are set without extensive search), as detailed below. This helps to reduce the possibility of unfair performance distortion due to excessive tuning.
  • SWAN-0 uses naive NS-iteration for whitening, disabled learning rate warmup, and use similar learning rates optimized for Adam. We fix the global learning rate to be the same as Adam, and fix $\alpha=1$. The only exception is the 1.3 B case. This is because we observe that the optimal learning rate of Adam under 1.3B becomes smaller than 0.001, hence we also reduce the learning rate on SWAN, where we used $\alpha=0.3$, resulting an effective learning rate of $0.0003$. To summarize the hyperparameter of SWAN-0 is set to be similar to Adam, without any tuning. This is to demonstrate the robustness of SWAN series optimizers and their capability to work out-of-the-box as a replacement for Adam.
  • SWAN, is the vanilla version of our method, in which we enabled learning rate warmup, and allowed the use of optimized learning rates that largely differ from Adam. We notice that SWAN allows larger learning rates than Adam. We use a global learning rate of 0.02, as well as the scaling factor $\alpha=0.05$. This is selected by simply searching the learning rate over a constraint grid $\{0.01,0.02,0.05\}$, and then setting $\alpha=0.05$ such that the effective learning rate is scaled back to 0.001. There is no guarantee that this heuristic rule is optimal; but we found that this usually does not make a run fail (e.g., with loss divergence).
  • Finally, SWAN, the most efficient version of SWAN that employs the proposed NSDS scheme for fast whitening (section 4.2). Similar to SWAN, we use the same global learning rate of 0.02, as well as the scaling factor $\alpha=0.05$ across all model sizes. We suspect with more careful tuning, its performance can be significantly improved; however, this is out of the scope of the paper.
The configurations of $\mathtt{GradWhitening}$ is discussed next.
Implementation of $\mathtt{GradWhitening}$
For $\mathtt{GradWhitening}$, before N-S iteration, we further normalize its input matrix by its Frobenius norm. This is applied in all our ablation studies as well, regardless of whether the gradient is processed by $\mathtt{GradNorm}$. For our proposed Newton-Schulz with Diagonal Substitution (NSDS) used in SWAN, we only run it for 2 steps across all model sizes. We found that using a step size $\beta\neq 0.5$ in the $\mathtt{GradWhitening}$ operator (Algorithm 2) can improve its convergence. We set $\beta=0.4$. We found that NSDS is generally robust to $\beta$ as long as it is not too large; and our specific choice of parameters usually already gives satisfactory performance in LLM pretraining. For the naive N-S iteration used in SWAN-0 and SWAN, we run 10 steps which is usually sufficient. We set $\beta=0.8$. The naive NS is run in FP32 precision, on top of BF16 gradients and weights; while the NSDS is run in BF16.
Computational Overhead.
[…]
However, note that whether the above analysis hold for large scale, distributed LLM training is still unclear. This is due to a few factors. First, the N-S iteration is after all a $\mathcal{O}(m^{3}$ complexity operation and might need to scale as model size increases. Second, the distributed computation of N-S steps needs extra non-trivial effort, which brings additional infrastructure challenges. This motivates us to propose the NSDS scheme, which enables Adam-level throughput without performance compromise.